OpenStudy (anonymous):

How to find the roots of t(x)= (x+5)^3+7??

OpenStudy (jdoe0001):

set the function to 0, that is \(\bf t(x)= (x+5)^3+7\implies 0= (x+5)^3+7\implies -7=(x+5)^3 \\ \quad \\ \Large \sqrt[3]{-7}=x+5\)

OpenStudy (anonymous):

youre awesome :-) @jdoe0001

OpenStudy (jdoe0001):

yw